# Exploiting the Layered Intrinsic Dimensionality of Deep Models for Practical Adversarial Training

### v1

@@ Appendix C Additional Results of Robustifying Sentiment Classifiers
Datasets. We evaluate SMAATon three datasets: AG-News Corpus (AGNEWS) [75], Internet Movie Database (IMDB) [37], and Yelp Review Polarity (YELP) [74]. The AGNEWS dataset contains over 120000 samples, each belonging to one of the four labels: World, Sports, Business, Sci/Tech. The IMDB dataset contains 50000 data samples of movie reviews with binary labels for negative and positive sentiments. The YELP dataset contains nearly 600000 samples of highly polar Yelp reviews with binary labels. However, due to limitations in computing resources, we only use a subset of 63000 samples of the YELP dataset. In addition, we randomly sample 10% of the training set for validation in all datasets. For testing, we use a subset of 1000 test samples from each dataset, following previous work practices. The AGNEWS dataset contains over 120k samples, categorized into four classes: World, Sports, Business, and Sci/Tech. The IMDB dataset consists of 50k movie reviews, each labeled with binary sentiments (positive or negative).
@@
Base model. We employed the $\text{BERT}_{\text{base-cased}}$ [15] and $\text{RoBERTa}_{\text{base-cased}}$ [34] models in our experiments. To conduct the evaluations, we utilize the fine-tuned models provided by TextAttack from HuggingFace for all datasets, except for the RoBERTa base model fine-tuned on YELP dataset. For the YELP dataset, we created a fine-tuned RoBERTa model for 2 epochs with a learning rate of $051e-$ 05 and a batch size of 32.
@@
Adversarial Attacks. which include the following constraints: (1) The maximum percentage of modified words is set to $0.3$ for AGNEWS, $0.1$ for IMDB and YELP datasets, respectively. (2) For word replacement, a maximum of $50$ candidates are considered for each word. (3) The semantic similarity, measured using the Universal Sentence Encoder [11], between the original input and the generated adversarial example must exceed $0.84$. PWWS uses word synonyms, TextFooler applies nearest neighbor search in counter-fitting embeddings [42], and BERT-Attack utilizes BERT masked language model to generate candidate words.
@@
Baselines. For input space adversarial training, we consider Adversarial Sparse Convex Combination (ASCC) [16] and Dirichlet Neighborhood Ensemble (DNE) [79]. These methods model the perturbation space as the convex hull of word synonyms. ASCC incorporates an entropy-based sparsity regularizer to capture word substitution geometry more effectively, while DNE expands the convex hull to encompass all synonyms of each word’s synonyms, combining Dirichlet sampling and adversarial training to enhance model robustness. In our investigation of embedding space adversarial training which recognized as the most impactful technique for enhancing generalization [33], we conduct a thorough analysis of FreeLB++ [32] (employs gradient-guided perturbations centered around the most susceptible data points). For certified defenses, we evaluate SAFER [71], TMD [40], and RSMI [40]. SAFER constructs a set of randomized inputs by performing random synonym substitutions and using the statistical properties of predicted labels to certify robustness. TMD employs infoGAN [12] to project adversarial examples to the data manifold in the last layer to address the manifold issue. RMSI combines these ideas by applying importance-based masking to tokens and leveraging randomized smoothing in each layer.
@@
Implementation details. To train the last layer of $f_{\theta}$ with adversarial samples, we create adversarial samples using 5-step PGD attacks. During training, we use epsilon values of $0.1$, $0.1$, and $0.8$ for the YELP, AGNEWS, and IMDB datasets, respectively, for the BERT models. For the RoBERTa models, we employ epsilon values of $0.1$, $0.6$, and $0.03$. All models are trained $10$ epochs with a learning rate of $0.1$.
Appendix D Additional Results of Language Understanding Benchmarks
To comprehensively evaluate SMAAT’s performance against a broader spectrum of textual adversarial attacks, we employ the GLUE and AdvGLUE benchmarks. The GLUE benchmark [63] is a comprehensive evaluation suite featuring seven diverse NLP tasks to assess model performance. The AdvGLUE benchmark [64] is an extension of GLUE, incorporating 17 distinct textual adversarial attacks, covering word-level transformations, sentence-level manipulations, and human-written AEs. This extension ensures a thorough evaluation encompassing various adversarial linguistic phenomena. For our assessment, we employ the evaluation sets of four datasets across three different tasks: Sentiment Analysis (SST-2), Duplicate Question Detection (QQP), and Natural Language Inference (QNLI, RTE).
In our evaluation, we compare SMAAT against standard BERT and RoBERTa models [^4] We use the fine-tuned models available from https://huggingface.co/JeremiahZ, as well as their FreeLB++ incorporated versions. In the case of SMAAT, we conducted a grid search for the learning rate, ranging from $0.1$ to $0.001$, and the $\epsilon$ value, ranging from $0.8$ to $0.01$, using 3-PGD steps As shown in Table D, SMAAT demonstrates a robustness improvement of 5.6% and 2.6% for BERT, and 12.1% and 2.5% for RoBERTa, compared to the standard and FreeLB++ models, respectively, while maintaining similar generalization performance.
@@ Appendix D Additional Results of Language Understanding Benchmarks
{NiceTabular}
c|ccc|ccc Dataset BERT   RoBERTa
@@
Appendix E Additional Results AT at the Last Layer
For vision models and decoder-based LLMs (dec-LLMs), the observed intrinsic dimension (ID) characteristics suggest that $l^{*}$ corresponds to the initial layer. This finding is in line with YOPO [72], which emphasizes the critical role of the initial layers for adversarial training (AT) in vision models. For the sake of completeness, we provide additional experiments on applying AT exclusively to the last layer of the vision and dec-LLMs models. For vision models, we train a Vision Transformer (ViT) model on the CIFAR-10 dataset. For dec-LLMs, we train the LLAMA-2-7b model with 8-bit quantized LoRA [14] on the SST-2 dataset using the SFTTrainer library from HuggingFace [1].
Figure 5: Robust accuracy obtained from standard training, AT in the input layer ($l_{1}$) and AT at the last layer. Results demonstrate that applying AT at the last layer doesn’t consistently lead to improved robustness. This is inline with our finding.
Our results from different types of models in Fig. 5 show that while applying adversarial training (AT) at the last layer improves robustness for enc-LLMs, it performs similarly to standard training for vision and dec-LLMs. These results motivate the ID analyses of SMAAT.
@@ Appendix F Additional Results of Robustifying Safety Filters In Decoder Based LLMs
The GCG attack is deliberately designed to bypass the safety alignment of LLMs by generating a response to a potentially harmful prompt through appending an adversarial suffix to the user’s input. This strategy leverages previous methodologies by using (1) an affirmative response tactic [67, 10] to direct the model’s output towards the attacker’s intended outcome (used for loss calculation) and (2) a mix of greedy and gradient-based discrete optimization [51] to pinpoint the most susceptible tokens. A key attribute of the GCG attack is its transferability, demonstrating that adversarial suffix designed for a specific prompt on one model can successfully affect a broad range of other models.
A suggested defense against the GCG attack involves incorporating a lightweight binary classifier model designed to identify harmful prompts [27]. It is important to note, however, that these classifiers can still be vulnerable to such attacks. To assess the effectiveness of SMAAT against the GCG attack, we conduct the attack by setting the suffix length to 20 tokens. The attack is generated over 50 iterations, with 100 trials per iteration. Rather than leveraging the attack’s transferability, we tailor an adversarial suffix for every individual prompt and model to enhance the attack’s impact. Our assessment is conducted using two datasets, AdvBench [81] and HH-RLHF [7], and involves three models: BERT, RoBERTa, and DistilBERT. AdvBench features 640 training prompts (320 harmful, 320 safe) and 240 test prompts (120 harmful, 120 safe), whereas HH-RLHF includes approximately 44K harmful and 44K helpful training prompts, with 2.3K test samples for each prompt category. In all cases, standard models are trained over 5 epochs with a learning rate of $1e^{-5}$. Table 6 details the training hyperparameters for SMAAT.

### v2

@@ Appendix D Additional Results of Robustifying Sentiment Classifiers
Datasets. We evaluate SMAATon three datasets: AG-News Corpus (AGNEWS) (Zhang et al., 2015), Internet Movie Database (IMDB) (Maas et al., 2011), and Yelp Review Polarity (YELP) (Zhang et al., 2015). The AGNEWS dataset contains over 120000 samples, each belonging to one of the four labels: World, Sports, Business, Sci/Tech. The IMDB dataset contains 50000 data samples of movie reviews with binary labels for negative and positive sentiments. The YELP dataset contains nearly 600000 samples of highly polar Yelp reviews with binary labels. However, due to limitations in computing resources, we only use a subset of 63000 samples of the YELP dataset. In addition, we randomly sample 10% of the training set for validation in all datasets. For testing, we use a subset of 1000 test samples from each dataset, following previous work practices. The AGNEWS dataset contains over 120k samples, categorized into four classes: World, Sports, Business, and Sci/Tech. The IMDB dataset consists of 50k movie reviews, each labeled with binary sentiments (positive or negative).
@@
Base model. We employed the $\text{BERT}_{\text{base-cased}}$ (Devlin et al., 2019) and $\text{RoBERTa}_{\text{base-cased}}$ (Liu et al., 2019) models in our experiments. To conduct the evaluations, we utilize the fine-tuned models provided by TextAttack from HuggingFace for all datasets, except for the RoBERTa base model fine-tuned on YELP dataset. For the YELP dataset, we created a fine-tuned RoBERTa model for 2 epochs with a learning rate of $051e-$ 05 and a batch size of 32.
@@
Adversarial Attacks. which include the following constraints: (1) The maximum percentage of modified words is set to $0.3$ for AGNEWS, $0.1$ for IMDB and YELP datasets, respectively. (2) For word replacement, a maximum of $50$ candidates are considered for each word. (3) The semantic similarity, measured using the Universal Sentence Encoder (Cer et al., 2018), between the original input and the generated adversarial example must exceed $0.84$. PWWS uses word synonyms, TextFooler applies nearest neighbor search in counter-fitting embeddings (Mrkšić et al., 2016), and BERT-Attack utilizes BERT masked language model to generate candidate words.
@@
Baselines. For input space adversarial training, we consider Adversarial Sparse Convex Combination (ASCC) (Dong et al., 2021) which model the perturbation space as the convex hull of word synonyms. ASCC incorporates an entropy-based sparsity regularizer to capture word substitution geometry more effectively. In our investigation of embedding space adversarial training which recognized as the most impactful technique for enhancing generalization (Li et al., 2021c), we conduct a thorough analysis of FreeLB++ (Li et al., 2021b) (employs gradient-guided perturbations centered around the most susceptible data points). For certified defenses, we evaluate SAFER (Ye et al., 2020), TMD (Minh & Luu, 2022), and RSMI (Minh & Luu, 2022). SAFER constructs a set of randomized inputs by performing random synonym substitutions and using the statistical properties of predicted labels to certify robustness. TMD employs infoGAN (Chen et al., 2016) to project adversarial examples to the data manifold in the last layer to address the manifold issue. RMSI combines these ideas by applying importance-based masking to tokens and leveraging randomized smoothing in each layer.
@@
Implementation details. To train the last layer of $f_{\theta}$ with adversarial samples, we create adversarial samples using 5-step PGD attacks. During training, we use epsilon values of $0.1$, $0.1$, and $0.8$ for the YELP, AGNEWS, and IMDB datasets, respectively, for the BERT models. For the RoBERTa models, we employ epsilon values of $0.1$, $0.6$, and $0.03$. All models are trained $10$ epochs with a learning rate of $0.1$. In our evaluation, we use a V100 GPU with 32 GB memory and 64 CPUs.
Appendix E Additional Results of Language Understanding Benchmarks
@@
Dataset BERT RoBERTa
@@
To comprehensively evaluate SMAAT’s performance against a broader spectrum of textual adversarial attacks, we employ the GLUE and AdvGLUE benchmarks. The GLUE benchmark (Wang et al., 2019a) is a comprehensive evaluation suite featuring seven diverse NLP tasks to assess model performance. The AdvGLUE benchmark (Wang et al., 2021) is an extension of GLUE, incorporating 17 distinct textual adversarial attacks, covering word-level transformations, sentence-level manipulations, and human-written AEs. This extension ensures a thorough evaluation encompassing various adversarial linguistic phenomena. For our assessment, we employ the evaluation sets of four datasets across three different tasks: Sentiment Analysis (SST-2), Duplicate Question Detection (QQP), and Natural Language Inference (QNLI, RTE).
In our evaluation, we compare SMAAT against standard BERT and RoBERTa models [^3] We use the fine-tuned models available from https://huggingface.co/JeremiahZ, as well as their FreeLB++ incorporated versions. In the case of SMAAT, we conducted a grid search for the learning rate, ranging from $0.1$ to $0.001$, and the $\epsilon$ value, ranging from $0.8$ to $0.01$, using 3-PGD steps As shown in Table 5, SMAAT demonstrates a robustness improvement of 5.6% and 2.6% for BERT, and 12.1% and 2.5% for RoBERTa, compared to the standard and FreeLB++ models, respectively, while maintaining similar generalization performance.
@@ Appendix F Additional Results of Robustifying Safety Filters In Decoder Based LLMs
The GCG attack is deliberately designed to bypass the safety alignment of LLMs by generating a response to a potentially harmful prompt through appending an adversarial suffix to the user’s input. This strategy leverages previous methodologies by using (1) an affirmative response tactic (Wei et al., 2023; Carlini et al., 2023) to direct the model’s output towards the attacker’s intended outcome (used for loss calculation) and (2) a mix of greedy and gradient-based discrete optimization (Shin et al., 2020) to pinpoint the most susceptible tokens. A key attribute of the GCG attack is its transferability, demonstrating that adversarial suffix designed for a specific prompt on one model can successfully affect a broad range of other models.
A suggested defense against the GCG attack involves incorporating a lightweight binary classifier model designed to identify harmful prompts (Kumar et al., 2023). It is important to note, however, that these classifiers can still be vulnerable to such attacks. To assess the effectiveness of SMAAT against the GCG attack, we conduct the attack by setting the suffix length to 20 tokens. The attack is generated over 50 iterations, with 100 trials per iteration. Rather than leveraging the attack’s transferability, we tailor an adversarial suffix for every individual prompt and model to enhance the attack’s impact. Our assessment is conducted using two datasets, AdvBench (Zou et al., 2023) and HH-RLHF (Bai et al., 2022), and involves three models: BERT, RoBERTa, and DistilBERT. AdvBench features 640 training prompts (320 harmful, 320 safe) and 240 test prompts (120 harmful, 120 safe), whereas HH-RLHF includes approximately 44K harmful and 44K helpful training prompts, with 2.3K test samples for each prompt category. In all cases, standard models are trained over 5 epochs with a learning rate of $1e^{-5}$. Table 6 details the training hyperparameters for SMAAT.